# A Hyperparameter Study for Quantum Kernel Methods

### v2

@@ IV-B4 Size of the reduced density matrices ($K$)
In Appendix E, we explore an alternative option for $\alpha_{k}$ in equations Eq. 4 and Eq. 5, which primarily impacts the influence of $K$ on the metrics. Furthermore, artificial labels for various $K$ lead to different observations depending on $\alpha_{k}$.
@@ IV-C Relabeled datasets
When the relabeling is done for $t=0.5$ the separation between the best classical and the best quantum accuracy is smaller compared to relabeling for $t=32$, which is displayed in Table III. This is expected because the gradient GD increases with $t$ as shown in Fig. 10. Similarly, higher values of $K$ also result in larger GD (see Section IV-B4) in the case of $t=32$ this results in a larger separation. However, when $t=0.5$, the GD is too small to significantly improve the quantum model’s ability to learn the artificial labels in comparison to the classical models. Small deviations in the separation can be explained by the process described in Section III-D. This process includes applying a regularization term ($\lambda Id$), utilizing a train-test-split, and incorporating additional randomness, which could explain any variance in the separation. Similar to Table III we discuss Table V in Appendix E, where we set $\alpha_{k}=1$ instead of $\alpha_{k}=1/N_{K}$ in Eq. 4 and Eq. 5.
@@ V Pipeline for investigating new datasets
Our experiments have shown the importance of the hyperparameter $t$, which is responsible for the data scaling and whose importance has been emphasized in [4, 24]. This raises an interesting question whether there is a good universal scaling factor for the Hamiltonian evolution feature map. To investigate this, we have found the values of $t$ that provide the best performance in terms of accuracy on average (over all other hyperparameter settings). Additionally, inner and distance kernels are considered separately as the distance kernel has an additional parameter $\gamma$, which exerts control over the scaling as well. As can be seen in Fig. 11, the datasets are scaled down to a similar range for both types of kernels separately. From here it follows, that scaling the data down to have standard deviation of approximately $0.49$ for the inner kernel and $0.22$ for the distance kernel would be a good initial guess.
@@ VI Discussion
The experiment demonstrated that the distance kernel outperforms the inner kernel when $\gamma$ is optimized, but when we average $\gamma$ over the range we investigated, the opposite is true (Fig. 5 and Fig. 16). The inner normalized and inner kernels have almost the same best performances when the inner is scaled to the same norm. Our results also indicate that the each hyperparameter has compatible importance for both the accuracy and the GD (Fig. 2). These importances, which are extracted from the GBDT model (Section III-E), match our intuition based in the standard deviations in the marginals and the formulas.
@@
The $\gamma$ hyperparameter is easy to optimize in terms of computational resources and because of its single extreme point. Since it can also make or break the distance model, it should be optimized. $\gamma$ also shows correlations with other parameters, most importantly with $t$. Since both $t$ and $\gamma$ regulate the bandwidth, they are interchangeable to a certain degree. Hence, the behavior of the distance kernel over the inner kernel when varying $t$ is more robust. This can be seen in Fig. 3(a), which also highlights the fact that the areas with good accuracy and the ones with high GD are rather distinct, especially regarding the dependence on $t$. The importance of the size of the RDM $K$ is relatively small for the setting we investigated (Fig. 2), but we assume that the importance of this parameter to increase with the number of features (qubit count) as the exponential concentration increases in the same way. Notably, the importance of $K$ also increases when we set $\alpha_{k}=1$ (Eq. 4 and Eq. 5), which is discussed in Appendix E. This choice however also leads to a stronger dependence of the bandwidth on $K$. $T$ has a relatively insignificant effect on the performance or the GD as already stated in [24]. This suggests that an exact evolution in the Hamiltonian evolution feature map is not necessary, meaning that the approximation one tries to reduce by Trotterization actually does not result in a worse embedding compared to the exact evolution. From the formula stated in Equation (Eq. 2), one cannot expect independence from the permutation of the features. In Appendix D we investigate this effect and determine that different permutations lead to different predictions, but that the impact is not significant.
@@
In supervised learning setup, we can consider on one hand the complexity of the features and on the other hand the complexity of the labels. That the complexity of the labels can completely change assumptions based on the features is the reason why the GD and the accuracy sometimes have different dependencies. It is also the reason why a relabeled dataset and the original one have little similarities in terms of accuracy despite having the same features. The GD, which solely depends on the features is therefore insufficient for indicating the performance as it was already clearly stated in [7], where this metric is only the first split point in the flowchart they suggested. Our experiments also empirically show that the GD cannot be used directly to tune hyperparameters, especially as an alternative to cross validation. Nevertheless, we found areas where the GD and performance behave similarly, oppositely, or independently to the accuracy. These implications might be of relevance in the search for real world datasets with an inductive bias that favors quantum KMs. For example, it might in general be better to have a higher $t$ with a lower $\gamma$ compared to a low $t$ with a high $\gamma$. Both possibilities can have the same accuracy, but the first one is closer to the area with a high GD as the second one based on Fig. 3(b) and therefore might be a better initial point. Choosing a larger value for $K$ notably increases the GD and slightly enhances performance. This, nonetheless, leads to Hilbert spaces with higher dimensions and convergence of the kernel to a fixed value [26]. RDM kernels might remedy this problem by contracting the space, which comes at an expense of the possible expressivity loss (as indicted by GD). $T$ and especially $C$ can be optimized rather independently of considerations of the potential for quantum advantage.
@@ VII Conclusion
To summarize our contributions, in this work we have performed an extensive study of the effects of different hyperparameter settings on the accuracy and the GD score. Our results suggest that even though the hyperparameters have comparable importance for both the accuracy and the GD, the optimal values peaks for the accuracy and GD do not align as well. The most important hyperparameter for both performance metrics is bandwidth $t$, which enforces the claims that were made in previous literature. This hyperparameter controls the scaling of the input data, and among the investigated datasets, the optimal $t$ corresponded to a scaling of the features in a similar range. The GD and the accuracy were observed to depend in an opposing manner on $t$. The distance kernel is more robust and performs better than the inner kernel thanks to $\gamma$ that can be easily optimized and has a similar effect as $t$. The two alternatives for providing the necessary kernel trace to compute the GD performed almost equally.
@@
In order to further increase the understanding of the GD we also suggest that more experiments with relabeled datasets, different embeddings and choices of RDM kernels are conducted. Further analyzing the spectral properties of the quantum kernel can shed light on the inductive bias of the model [10, 25]. Another interesting future work in context of the quantum Hamiltonian evolution feature map would be to include noise in such a broad hyperparameter analysis, which is ideally done with real hardware runs. Increasing the number of qubits may provide new insights, particularly regarding the role of $K$.
@@ Acknowledgements
Appendix A Analysing different kernel functions
@@
Appendix B Additional details of empirical study
@@
Appendix C Effects of changing dataset preprocessing steps
@@
Appendix D Effects of feature permutations
@@
Appendix E Effects of $\alpha_{K}$ parameter

### v3

@@ IV-B3 Gamma ($\gamma$)
Figure 5: Dependence of the test accuracy on the choice of basis for the kernel function. For this comparison $\gamma$ is optimized and not averaged. The specific form of a distance- or inner based kernel function is given in Section II-C. The dashed line indicates the accuracy that would be achieved by random guessing.
@@ IV-B4 Size of the reduced density matrices ($K$)
In Appendix F, we explore an alternative option for $\alpha_{k}$ in equations Eq. 4 and Eq. 5, which primarily impacts the influence of $K$ on the metrics. Furthermore, artificial labels for various $K$ lead to different observations depending on $\alpha_{k}$.
@@ IV-C Relabeled datasets
When the relabeling is done for $t=0.5$ the separation between the best classical and the best quantum accuracy is smaller compared to relabeling for $t=32$, which is displayed in Table III. This is expected because the gradient GD increases with $t$ as shown in Fig. 10. Similarly, higher values of $K$ also result in larger GD (see Section IV-B4) in the case of $t=32$ this results in a larger separation. However, when $t=0.5$, the GD is too small to significantly improve the quantum model’s ability to learn the artificial labels in comparison to the classical models. Small deviations in the separation can be explained by the process described in Section III-D. This process includes applying a regularization term ($\lambda Id$), utilizing a train-test-split, and incorporating additional randomness, which could explain any variance in the separation. Similar to Table III we discuss Table V in Appendix F, where we set $\alpha_{k}=1$ instead of $\alpha_{k}=1/N_{K}$ in Eq. 4 and Eq. 5.
@@ V Pipeline for investigating new datasets
Our experiments have shown the importance of the hyperparameter $t$, which is responsible for the data scaling and whose importance has been emphasized in [4, 24]. This raises an interesting question whether there is a good universal scaling factor for the Hamiltonian evolution feature map. To investigate this, we have found the values of $t$ that provide the best performance in terms of accuracy on average (over all other hyperparameter settings). Additionally, inner and distance kernels are considered separately as the distance kernel has an additional parameter $\gamma$, which exerts control over the scaling as well. As can be seen in Fig. 11, the datasets are scaled down to a similar range for both types of kernels separately. From here, it follows, that scaling the data down to have standard deviation of approximately $0.49$ for the inner kernel and $0.22$ for the distance kernel would be a good initial guess.
@@ VI Discussion
The experiment demonstrated that the distance kernel outperforms the inner kernel when $\gamma$ is optimized, but when we average $\gamma$ over the range we investigated, the opposite is true (Fig. 5 and Fig. 17). Our results also indicate that the each hyperparameter has compatible importance for both the accuracy and the GD (Fig. 2). These importances, which are extracted from the GBDT model (Section III-E), match our intuition based in the standard deviations in the marginals and the formulas.
@@
The $\gamma$ hyperparameter is easy to optimize in terms of computational resources and because of its single extreme point. Since it can also make or break the distance model, it should be optimized. $\gamma$ also shows correlations with other parameters, most importantly with $t$. Since both $t$ and $\gamma$ regulate the bandwidth, they are interchangeable to a certain degree. Hence, the behavior of the distance kernel over the inner kernel when varying $t$ is more robust. This can be seen in Fig. 4(a), which also highlights the fact that the areas with good accuracy and the ones with high GD are rather distinct, especially regarding the dependence on $t$. The importance of the size of the RDM $K$ is relatively small for the setting we investigated (Fig. 2), but we assume that the importance of this parameter to increase with the number of features (qubit count) as the exponential concentration increases in the same way. Notably, the importance of $K$ also increases when we set $\alpha_{k}=1$ (Eq. 4 and Eq. 5), which is discussed in Appendix F. This choice however also leads to a stronger dependence of the bandwidth on $K$. $T$ has a relatively insignificant effect on the performance or the GD as already stated in [24]. This suggests that an exact evolution in the Hamiltonian evolution feature map is not necessary, meaning that the approximation one tries to reduce by Trotterization actually does not result in a worse embedding compared to the exact evolution. From the formula stated in Equation (Eq. 2), one cannot expect independence from the permutation of the features. In Appendix E we investigate this effect and determine that different permutations lead to different predictions, but that the impact is not significant.
@@
In a supervised learning setup, we can consider on one hand the complexity of the features and on the other hand the complexity of the labels. The complexity of the labels can completely change assumptions based on the features is the reason why the GD and the accuracy sometimes have different dependencies. It is also the reason why a relabeled dataset and the original one have little similarities in terms of accuracy despite having the same features. Understanding the intricate relationship between the characteristics of the features, complexity of the labels, GD metric and empirical difference between quantum and classical learners can help us in the search for real world datasets that favor the inductive bias of quantum KMs. The results from this study provide us with initial hints. For example, it might in general be better to have a higher $t$ with a lower $\gamma$ compared to a low $t$ with a high $\gamma$. Both possibilities can have the same accuracy, but the first one is closer to the area with a high GD as the second one based on Fig. 4(b) and therefore might be a better initial point. Choosing a larger value for $K$ notably increases the GD and slightly enhances performance. This, nonetheless, leads to Hilbert spaces with higher dimensions and convergence of the kernel to a fixed value [26]. RDM kernels might remedy this problem by contracting the space, which comes at an expense of the possible expressivity loss (as indicated by GD). Further research is required to understand use-cases that might particularly benefit from the deployment of quantum models.
@@ VII Conclusion
To summarize our contributions, in this work we have performed an extensive study of the effects of different hyperparameter settings on the accuracy and the GD score. Our results suggest that even though the hyperparameters have comparable importance for both the accuracy and the GD, the optimal values peaks for the accuracy and GD sometimes differ. This study provides some clues towards understanding the difference between classical and quantum kernel methods. The most important hyperparameter for both performance metrics is bandwidth $t$, which enforces the claims that were made in previous literature. This hyperparameter controls the scaling of the input data, and among the investigated datasets, the optimal $t$ corresponded to a scaling of the features in a similar range. The GD and the accuracy were observed to depend in an opposing manner on $t$. The distance kernel is more robust and performs better than the inner kernel thanks to $\gamma$ that can be easily optimized and has a similar effect as $t$. The two alternatives for providing the necessary kernel trace to compute the GD performed almost equally.
@@
In order to further increase the understanding of the GD we also suggest that more experiments with relabeled datasets, different embeddings and choices of RDM kernels are conducted. Further analyzing the spectral properties of the quantum kernel can shed light on the inductive bias of the model [10, 25]. Another interesting future work in context of the quantum Hamiltonian evolution feature map would be to include noise in such a broad hyperparameter analysis, which is ideally done with real hardware runs. Both noisy runs and increasing the number of qubits may provide new insights, particularly regarding the role of $K$, as exponential concentration [26] would become more pronounced. This hyperparameter governs the size of the projection that helps counteract the effects of the exponential concentration, and, hence, its importance could rise once we consider scenarios that are more affected by this phenomenon.
Given the relationship of QKMs to other QML approaches [22], an interesting further research direction would be to investigate how our findings translate to other models. For example, hyperparameters responsible for scaling of the input, such as $t$ and $\gamma$, have been identified to have a dominating influence while their optimal values put the data ranges (see Fig. 11) way below the industrial standard. This begs the question whether other QML models, such as Quantum Neural Networks, might benefit from a change in the standard data preprocessing procedure to fit the preprocessed data into the identified ranges.
@@ Acknowledgements
Appendix A Inner normalized kernel function
Additionally to the kernels considered in the main text, we tested an adaptation of the inner product based kernel. The motivation is that the kernel entries can become relatively small for large dimensions and highly entangled states. On the main diagonal the values are related to the purity of a quantum state $Tr(\rho^{2})$, which can become $\frac{1}{d}$ where $d$ is the dimension of the Hilbert space. The normalized inner product based kernel is defined as
and therefore remains a valid kernel, because the same feature vectors as in Nakaji et al. [19] can be found for the redefined density matrices (DMs). This normalization affects all entries in the gram matrix, in particular the main diagonal now consists of just ones, as it is common for kernels in general. Important to note is that this kernel is only meaningful in a simulation environment.
When comparing methods inner and inner normalized we notice almost no difference in performance (as seen in Fig. 13) or GD (as seen in later sections). This implies that this adaptation has almost no influence on classification of the SVM.
Appendix B Analysing different kernel functions
@@
Appendix C Additional details of empirical study
@@
Appendix D Effects of changing dataset preprocessing steps
@@
Appendix E Effects of feature permutations
@@
Appendix F Effects of $\alpha_{K}$ parameter